# TempMe: Video Temporal Token Merging for Efficient Text-Video Retrieval

### v1

@@ Abstract
Most text-video retrieval methods utilize the text-image pre-trained CLIP as a backbone, incorporating complex modules that result in high computational overhead. As a result, many studies focus on efficient fine-tuning. The primary challenge in efficient adaption arises from the inherent differences between image and video modalities. Each sampled video frame must be processed by the image encoder independently, which increases complexity and complicates practical deployment. Although existing efficient methods fine-tune with small trainable parameters, they still incur high inference costs due to the large token number. In this work, we argue that temporal redundancy significantly contributes to the model’s high complexity due to the repeated information in consecutive frames. Existing token compression methods for image models fail to solve the unique challenges, as they overlook temporal redundancy across frames. To tackle these problems, we propose Temporal Token Merging (TempMe) to reduce temporal redundancy. Specifically, we introduce a progressive multi-granularity framework. By gradually combining neighboring clips, we merge temporal tokens across different frames and learn video-level features, leading to lower complexity and better performance. Extensive experiments validate the superiority of our TempMe. Compared to previous efficient text-video retrieval methods, TempMe significantly reduces output tokens by $\mathbf{95\%}$ and GFLOPs by $\mathbf{51\%}$, while achieving a $\mathbf{1.8\times}$ speedup and a $\mathbf{4.4\%}$ R-Sum improvement. Additionally, TempMe exhibits robust generalization capabilities by integrating effectively with both efficient and full fine-tuning methods. With full fine-tuning, TempMe achieves a significant $\mathbf{7.9\%}$ R-Sum improvement, trains $\mathbf{1.57\times}$ faster, and utilizes $\mathbf{75.2\%}$ GPU memory usage. Our code will be released.
00footnotetext: Equal contribution. Corresponding authors.
@@ 1 Introduction
In the domain of video-language understanding, text-video retrieval is a crucial task focused on matching videos that correspond to specific query texts or vice versa. Given the powerful representation capacity of large-scale text-image pre-training [10, 40, 48], recent studies [36, 47, 39, 20] have increasingly focused on full fine-tuning of large foundational models for text-video retrieval. Specifically, leveraging the pre-trained CLIP [40] as the backbone, present full fine-tuning approaches incorporate cumbersome modules to extract video and text representations. However, the slow inference speed of these approaches severely limits their real-world applications. For example, the training process of CLIP4Clip [36] with ViT-B/16 requires 70.1GB GPU memory usage and takes 6.5 hours. Therefore, we focus on efficient fine-tuning text-video retrieval [23, 18, 53] to lower computational expenses.
@@
There are substantial challenges in efficiently adapting a text-image pre-trained model for text-video retrieval, mainly due to the inherent differences between image and video modalities. Since videos consist of numerous frame images, each sampled frame in a video needs to pass through the pre-trained image encoder, leading to high complexity. The shift from processing a single image to handling multiple sampled frames dramatically raises the number of patch tokens fed into the model, complicating the extraction of video representations. As a result, this adaption suffers from high computational overhead, which makes practical deployment difficult. As shown in Figure 1b, existing efficient text-video retrieval methods like VoP [18] and DGL [53] freeze the pre-trained backbone and introduce trainable prompts to model video data. They treat video input as a sequence of multiple sampled frames, acquiring considerable accuracy at the expense of inference efficiency.
@@
Figure 1: (a) An example illustrates the large temporal redundancy between adjacent frames. Identical subjects are highlighted in the same color. (b) Current methods treat video input as a sequence of multiple image frames, causing high complexity due to the large number of tokens. (c) In contrast, our TempMe reduces temporal redundancy by progressively merging redundant tokens in adjacent video clips. (d) With ViT-B/16 as the backbone on MSRVTT, our TempMe reaches state-of-the-art performance with minimal computational overhead. R-Sum is the sum of R@1, R@5, and R@10.
@@
When employing the pre-trained image encoder for video inputs, we argue that temporal redundancy plays a major role in the model’s high complexity. The sequential sampled frames within a video capture a continuous temporal progression, illustrating subject evolution, interaction, and transition over time. As shown in Figure 1a, the video depicts a cat and a monkey in a fight, showing the same subjects across adjacent frames while maintaining a repetitive background. This temporal progression introduces substantial redundancy due to repeated information present in consecutive frames, contributing to the high complexity caused by repeated model calculations.
@@
A recently popular technique, token compression [41, 26, 6, 4], might be a fascinating solution, which has achieved huge success in accelerating image models. Classically, token compression methods aim to eliminate spatial redundancy and merge similar tokens in the image to improve efficiency. However, they fail to resolve the unique challenges imposed by video modality. Specifically, they overlook the temporal redundancy across different frames, leading to inferior performance. Currently, there is no dedicated compression algorithm designed for text-video retrieval models, and the efficiency issue urgently needs to be addressed.
@@
In this paper, we introduce Temporal Token Merging (TempMe) for efficient text-video retrieval, a novel method designed to minimize inherent temporal redundancy in video content. Our proposed TempMe effectively merges a large number of similar tokens across frames, significantly reducing computational overhead. Particularly, our method adopts a progressive multi-granularity framework, enabling the efficient merging of redundant tokens, as visualized in Figure 1c. Videos can be viewed as aggregations of clips with varying temporal granular levels, from individual images at the micro-level to the entire video at the macro-level. Through gradually combining neighboring clips, the total number of tokens is dramatically decreased, while the features evolve from fine-grained image-level features to holistic video-level features, leading to lower complexity and better performance.
@@
To validate the advantages of our TempMe, extensive experiments are conducted on four benchmark datasets, MSRVTT [49], ActivityNet [27], DiDeMo [1], and LSMDC [43]. Experimental results consistently demonstrate that our TempMe offers a leading balance between speed and accuracy, outperforming other efficient fine-tuning methods. Figure 1d shows the performance comparison of ViT-B/16 on MSR-VTT. Compared to VoP [18] and DGL [53], TempMe significantly reduces output tokens by $\mathbf{95\%}$ and GFLOPs by $\mathbf{51\%}$, while achieving a $\mathbf{1.8\times}$ speedup and a $\mathbf{4.4\%}$ R-Sum improvement. Moreover, TempMe can be seamlessly integrated with various efficient fine-tuning (see Appendix C) and full fine-tuning (see Table 5), which demonstrates its robust generalization capabilities. When fully fine-tuning with ViT-B/16, TempMe achieves a significant $\mathbf{7.9\%}$ R-Sum improvement, trains $\mathbf{1.57\times}$ faster, and utilizes $\mathbf{75.2\%}$ GPU memory usage.
@@
Our major contributions are summarized as follows: (1) Our study reveals that temporal redundancy in video content causes excessive computational demands when adapting pre-trained text-image models for text-video retrieval. The image encoder redundantly processes overlapping information across frames, resulting in increased model complexity. (2) To overcome such limitations, we propose TempMe for efficient text-video retrieval to reduce temporal redundancy. A progressive multi-granularity framework is proposed to merge redundant tokens in adjacent video clips step by step, simplifying model complexity while extracting unified and discriminative video features. (3) Experimental evaluation on four benchmark datasets demonstrates that our TempMe outperforms current state-of-the-art methods. Extensive experiments indicate that our TempMe effectively reduces model complexity while achieving superior performance.
@@ 2 Related Works
Due to the great success of language-vision pre-training [46, 51, 8, 31, 10, 40, 48, 30, 29] in various downstream cross-modal tasks [2, 50, 24, 52, 44], numerous studies [36, 13, 33, 47, 14, 21, 39, 20, 3, 28] have achieved remarkable results by utilizing CLIP’s [40] knowledge for text-video retrieval. These approaches typically employ CLIP’s encoders as the backbone and use complex similarity calculations. Specifically, CLIP4Clip [36] introduces several video aggregation schemes to obtain video features. However, these full fine-tuning methods with their customized modules require significant memory and computational resources. Therefore, our work aims to explore efficient fine-tuning methods to lower computational overhead.
Mainstream parameter-efficient fine-tuning approaches in the image domain, such as Prompt [25, 19, 57, 56], LoRA [17, 55], and Adapter [16, 15, 7, 55], freeze pre-trained model parameters while adding extra tunable parameters to minimize storage demands. For parameter-efficient text-video retrieval, VoP [18] and DGL [53] introduce extra modules to generate prompt tokens and capture global video information. Despite their good performance with small tunable parameters, these methods face high inference costs due to the significant token number in videos.
@@
Token compression methods [42, 32, 45, 34, 41, 26, 6, 4] aim to reduce computational burden by reducing tokens while preserving the essential information. Some works [38, 54, 11] focus on dynamically pruning non-essential tokens, leading to varying token numbers across images and thus impacting speed efficiency. Alternatively, some works [37, 4, 6] focus on merging tokens. ToMe [4] reduces tokens in a transformer gradually by merging a fixed number of tokens in each block. DiffRate [6] performs token pruning and merging simultaneously. DiffRate and ToMe can be combined with parameter-efficient fine-tuning methods to reduce tokens in text-video retrieval. Although ToMe has been applied to video classification by simultaneously processing all frame tokens in Spatiotemporal MAE [12], this implementation is incompatible with CLIP’s image encoder for text-video retrieval. When processing each frame via CLIP’s image encoder, ToMe effectively reduces spatial redundancy by merging similar tokens within a single frame. However, it does not address the temporal redundancy across frames, which greatly contributes to computational overhead.
Inspired by ToMe, we merge similar tokens within individual frames during the first image merging stage. Then, we propose an innovative clip merging stage for cross-clip and intra-clip merging, which significantly reduces temporal tokens and enhances the learning of video-level features.
@@ 3 Methodology
In this work, we concentrate on efficiently adapting the pre-trained text-image model to text-video retrieval. Firstly, preliminaries are presented in Section 3.1. Subsequently, the implementation details of our TempMe are elaborated in Section 3.2. Finally, complexity analysis is discussed in Section 3.3
@@ 3.1 Preliminaries
The task of text-video retrieval involves leveraging textual queries to accurately retrieve related videos, as well as utilizing video queries to find associated textual descriptions. The primary goal of text-video retrieval is to bridge the semantic gap between textual description and video content by learning a similarity function $s(t,v)$. Here, $t$ denotes a text, and $v$ represents a sequence of frame images $\left\{I_{i}\right\}_{i=1}^{F}$, where $F$ is the number of sampled frames in time. We employ LoRA [17] to adapt CLIP [40] for text-video retrieval tasks. The trainable matrices of LoRA can be merged, introducing no extra inference latency. Further details are provided in Appendix A.
@@ 3.2 Temporal Token Merging
For efficient text-video retrieval, we freeze the pre-trained CLIP and merely train LoRA in both the image and text encoders. In this subsection, we focus on merging temporal tokens across multiple frames to minimize temporal redundancy and extract more comprehensive video features.
@@
In text-video retrieval, the significant impact of temporal redundancy on model complexity has not been adequately explored. To mitigate temporal redundancy, we proposed the Progressive Multi-Granularity (PMG) framework that efficiently consolidates similar temporal tokens across different frames. In our PMG framework (see Figure 2), we introduce a two-stage merging approach: image merging and clip merging. Initially, ImgMe block independently encodes each single frame, merging similar tokens within the same frame and leveraging pre-trained image knowledge to capture low-level spatial details. Next, ClipMe block (see Figure 3) is proposed to aggregate short-frame clips into extended-frame clips, merging temporal tokens across different frames and learning unified and discriminative video features.
Specifically, each frame image $I$ can be regarded as a discrete video clip lasting only one frame $v^{f=1}$. Concatenating these consecutive clips yields an extended clip with increased temporal length. In the toy example presented in Figure 2, 8 individual frames $\left\{v^{f=1}_{i}\right\}_{i=1}^{8}$ are aggregated into four clips, each containing with two frames $\left\{v^{f=2}_{i}\right\}_{i=1}^{4}$. These clips are then aggregated to form a complete video clip with all 8 frames $\left\{v^{f=8}_{i}\right\}_{i=1}^{1}$. As clips are progressively aggregated, we merge a significant number of similar tokens that arise from overlapping information across frames. Simultaneously, comprehensive video information is captured through the attention mechanism that processes tokens across different frames.
@@
Inspired by ToMe [4], the image merging step merges tokens to reduce by $r$ per layer, which is employed between the MHSA and MLP branches of each ImgMe Block. Following Bipartite Soft Matching [4], we first partition the tokens into two sets $\mathbb{A}$ and $\mathbb{B}$ by alternating. For each token in set $\mathbb{A}$, we find its most similar token in Set $\mathbb{B}$ using cosine similarity. The top $r$ most similar pairs are then merged by averaging their features. Given an input $\mathbb{R}^{1\times N\times D}$, the image merging stage with $K$ ImgMe Block outputs $\mathbb{R}^{1\times(N-rK)\times D}$, where $N$ denotes the number of tokens, and $D$ denotes the dimension of the token feature.
To prevent loss of image information, $r$ is set to a low value: $r=2$ for ViT-B/32 and $r=10$ for ViT-B/16. Nevertheless, the model complexity remains high because of the processing of multiple frames. Therefore, we aim to reduce temporal tokens to further mitigate the model complexity.
@@
Instead of ImgMe Block which compresses tokens within each single frame, we propose ClipMe Block to process multi-frame clips as illustrated in Figure 3. In ClipMe Block, we introduce two novel steps: cross-clip merging and intra-clip merging. During cross-clip merging, adjacent clips are aggregated, which significantly reduces the number of temporal tokens and generates a new clip. The intra-clip merging step further compresses the tokens within the newly formed clip.
In cross-clip merging, we aim to combine tokens with their most similar tokens from the same subject, whether these matches occur within a single frame or span across multiple frames. Initially, all tokens from the input $f$ clips $\mathbb{R}^{f\times N\times D}$ are alternately divided into Set $\mathbb{A}$ or Set $\mathbb{B}$. This alternating assignment scheme is based on the observation that the adjacent tokens often depict the same subject, thus their features tend to be similar. Moreover, this assignment ensures that tokens representing the same subject across various frames are distributed between both sets. Considering the high temporal redundancy observed in videos, we select a large subset, $fN\times(1-R_{c})$, of the most similar token pairs for merging. Before merging, the trainable clip positional embeddings are added to facilitate temporal understanding in MSHA. Subsequently, these merged tokens, along with the unmerged tokens, constitute a new clip $\mathbb{R}^{1\times(fN\times R_{c})\times D}$, which is then fed into MHSA.
During intra-clip merging, the tokens in the newly formed clip, improved by MHSA for better temporal and spatial details, are further reduced. In contrast to the image merging step that merges a small number $r$ of tokens within an individual frame, the intra-clip merging step merges a larger proportion of tokens within a clip containing multiple frames, compressing $\mathbb{R}^{1\times N\times D}$ into $\mathbb{R}^{1\times(N\times R_{I})\times D}$.
Table 1: Complexity comparisons. We report the R-Sum metric for 12 sampled frames on MSRVTT and 64 sampled frames on ActivityNet, respectively. The best metric of all methods is highlighted in bold.
@@
Method GFLOPs # Tokens R-Sum GFLOPs # Tokens R-Sum GFLOPs # Tokens R-Sum
LoRA [17] 53.0 (100%) $12\times 50$ (100%) 193.0 276.7 (100%) $64\times~{}~{}50$ (100%) 191.5 211.3 (100%) $12\times 197$ (100%) 201.4
DiffRate [6] 36.8 (  69%) $12\times 20$ (  40%) 189.9 190.1 (  69%) $64\times~{}~{}20$ (  40%) 188.7 138.5 (  66%) $12\times~{}~{}49$ (  25%) 202.4
ToMe [4] 40.2 (  76%) $12\times 26$ (  52%) 191.4 208.5 (  75%) $64\times~{}~{}26$ (  52%) 189.6 144.4 (  68%) $12\times~{}~{}77$ (  39%) 200.6
TempMe 34.8 (  65%) $\mathbf{1\times 97}$ (  16%) 198.6 180.3 (  65%) $\mathbf{1\times 500}$ (  16%) 205.6 121.4 (  57%) $\mathbf{1\times 127}$ (    5%) 206.7

### v2

@@ Abstract
Most text-video retrieval methods utilize the text-image pre-trained models like CLIP as a backbone. These methods process each sampled frame independently by the image encoder, resulting in high computational overhead and limiting practical deployment. Addressing this, we focus on efficient text-video retrieval by tackling two key challenges: 1. From the perspective of trainable parameters, current parameter-efficient fine-tuning methods incur high inference costs; 2. From the perspective of model complexity, current token compression methods are mainly designed for images to reduce spatial redundancy but overlook temporal redundancy in consecutive frames of a video. To tackle these challenges, we propose Temporal Token Merging (TempMe), a parameter-efficient and training-inference efficient text-video retrieval architecture that minimizes trainable parameters and model complexity. Specifically, we introduce a progressive multi-granularity framework. By gradually combining neighboring clips, we reduce spatio-temporal redundancy and enhance temporal modeling across different frames, leading to improved efficiency and performance. Extensive experiments validate the superiority of our TempMe. Compared to previous parameter-efficient text-video retrieval methods, TempMe achieves superior performance with just 0.50M trainable parameters. It significantly reduces output tokens by $\mathbf{95\%}$ and GFLOPs by $\mathbf{51\%}$, while achieving a $\mathbf{1.8\times}$ speedup and a $\mathbf{4.4\%}$ R-Sum improvement. With full fine-tuning, TempMe achieves a significant $\mathbf{7.9\%}$ R-Sum improvement, trains $\mathbf{1.57\times}$ faster, and utilizes $\mathbf{75.2\%}$ GPU memory usage. The code is available at https://github.com/LunarShen/TempMe.
@@ 1 Introduction
In the domain of video-language understanding, text-video retrieval is a crucial task focused on matching videos that correspond to specific query texts or vice versa. Given the powerful capacity of large-scale text-image pre-training Fan et al. (2024); Radford et al. (2021); Xu et al. (2024a), recent studies Luo et al. (2022); Wang et al. (2023); Pei et al. (2023); Jin et al. (2022) have increasingly focused on full fine-tuning of CLIP Radford et al. (2021) for text-video retrieval. Specifically, these methods incorporate cumbersome modules to extract video and text representations. However, the slow inference speed of these methods severely limits their real-world applications. For example, the training process of CLIP4Clip Luo et al. (2022) with CLIP-ViT-B/16 requires 70.1GB GPU memory usage and takes 6.5 hours. Therefore, we focus on efficient fine-tuning text-video retrieval Ju et al. (2022); Huang et al. (2023); Yang et al. (2024) to lower computational expenses, which freeze the pre-trained backbone and introduce minimal trainable parameters to model video data.
@@
Efficient adaptation of text-image pre-trained models for text-video retrieval remains challenging due to the inherent differences between image and video modalities. The shift from processing a single image to handling multiple sampled frames dramatically raises the number of patch tokens fed into the model, complicating the extraction of video representations. As a result, this adaption suffers from high computational overhead, which makes practical deployment difficult. As shown in Figure 1b, existing parameter-efficient text-video retrieval methods  Huang et al. (2023); Yang et al. (2024); Jin et al. (2024); Cao et al. (2024) treat video input as a sequence of multiple sampled frames, achieving competitive accuracy but at the cost of inference efficiency.
@@
Figure 1: (a) An example illustrates the large temporal redundancy between adjacent frames. Identical subjects are highlighted in the same color. (b) Current methods treat video input as a sequence of multiple sampled frames, causing high complexity due to the large number of tokens. (c) In contrast, our TempMe reduces temporal redundancy by progressively merging redundant tokens in adjacent video clips. (d) With CLIP-ViT-B/16 on MSRVTT, our TempMe reaches state-of-the-art performance with minimal computational overhead. R-Sum is the sum of R@1, R@5, and R@10.
@@
Motivated by these limitations, we argue that temporal redundancy plays a major role in the model’s high complexity when employing the pre-trained image encoder for video inputs. The sequentially sampled frames in a video capture a continuous temporal progression, illustrating subject evolution, interaction, and transition over time. In Figure 1a, the video shows a cat and a monkey fighting, showing the same subjects in adjacent frames against a repetitive background. This temporal progression introduces substantial redundancy due to repeated information in consecutive frames, contributing to the high complexity resulting from repeated model calculations.
@@
To mitigate computational complexity, a recently popular technique, token compression Rao et al. (2021); Kong et al. (2022); Chen et al. (2023a); Bolya et al. (2022), might be a fascinating solution, which has achieved huge success in accelerating image models. Classically, token compression methods aim to merge redundant spatial tokens in the image to improve efficiency. However, they fail to resolve the unique challenges imposed by video modality. Specifically, they overlook the temporal redundancy across different frames, leading to inferior performance. Currently, there is no dedicated compression algorithm designed for text-video retrieval models, and the efficiency issue urgently needs to be addressed.
@@
In conclusion, existing methods struggle to achieve efficient text-video retrieval: (1) Parameter-efficient methods still lead to high inference costs. (2) Token compression methods fail to address temporal redundancy. In this paper, we propose Temporal Token Merging (TempMe), which effectively integrates parameter-efficient tuning with token compression to overcome these limitations. Particularly, our method adopts a progressive multi-granularity framework, enabling the efficient merging of redundant temporal tokens, as visualized in Figure 1c. Videos can be viewed as aggregations of clips with varying temporal granular levels, from individual images at the micro-level to the entire video at the macro-level. Through gradually combining neighboring clips, the total number of tokens is dramatically decreased, while the features evolve from fine-grained image-level features to holistic video-level features, leading to lower complexity and better performance.
@@
To validate the advantages of our TempMe, extensive experiments are conducted on four benchmark datasets, MSRVTT Xu et al. (2016), ActivityNet Krishna et al. (2017), DiDeMo Anne Hendricks et al. (2017), and LSMDC Rohrbach et al. (2015). Experimental results consistently demonstrate that our TempMe offers a leading balance between speed and accuracy, outperforming other efficient fine-tuning methods. Figure 1d shows the performance comparison of CLIP-ViT-B/16 on MSR-VTT. Compared to VoP Huang et al. (2023) and DGL Yang et al. (2024), TempMe significantly reduces output tokens by $\mathbf{95\%}$ and GFLOPs by $\mathbf{51\%}$, while achieving a $\mathbf{1.8\times}$ speedup and a $\mathbf{4.4\%}$ R-Sum improvement. Moreover, TempMe can be seamlessly integrated with various parameter-efficient fine-tuning (see Appendix C.2) and full fine-tuning (see Table 5), which demonstrates its robust generalization capabilities. When fully fine-tuning with CLIP-ViT-B/16, TempMe achieves a significant $\mathbf{7.9\%}$ R-Sum improvement, trains $\mathbf{1.57\times}$ faster, and utilizes $\mathbf{75.2\%}$ GPU memory usage.
@@
Our major contributions are summarized as follows: (1) Our study reveals that temporal redundancy in video content causes excessive computational demands when adapting pre-trained text-image models for text-video retrieval. Existing parameter-efficient methods incur high inference costs, while current token compression methods fail to address temporal redundancy. (2) To overcome such limitations, we propose TempMe, a parameter- and inference-efficient text-video retrieval architecture that merges redundant temporal tokens in adjacent video clips step by step, simplifying model complexity while extracting unified and discriminative video features. (3) Results on four benchmarks reveal that TempMe outperforms current SOTA methods, including both parameter-efficient video-text retrieval and compression methods. Extensive experiments indicate that our TempMe effectively reduces model complexity while achieving superior performance.
@@ 2 Related Works
Due to the great success of language-vision pre-training Wang et al. (2022b); Xue et al. (2022); Chen et al. (2023b); Li et al. (2023); Fan et al. (2024); Radford et al. (2021); Xu et al. (2024a); Li et al. (2021; 2022) in various downstream cross-modal tasks Antol et al. (2015); Xu et al. (2015); Karpathy & Fei-Fei (2015); Shen et al. (2024), numerous studies Luo et al. (2022); Gorti et al. (2022); Liu et al. (2022); Wang et al. (2023); Guan et al. (2023); Jin et al. (2023b); Pei et al. (2023); Jin et al. (2022); Shen et al. (2025); Lei et al. (2021) have achieved remarkable results by utilizing CLIP’s Radford et al. (2021) knowledge for text-video retrieval. These approaches typically employ CLIP’s encoders as the backbone and use complex similarity calculations. Specifically, CLIP4Clip Luo et al. (2022) introduces several video aggregation schemes to obtain video features. Cap4Video Wu et al. (2023) generates associated captions to improve performance. HBI Jin et al. (2023a) formulates video and text as players in a cooperative game. However, these full fine-tuning methods with their customized modules require significant memory and computational resources. Therefore, our work aims to explore efficient fine-tuning methods to lower computational overhead.
Mainstream parameter-efficient fine-tuning methods in the image domain, such as Prompt Khattak et al. (2023); Jia et al. (2022); Zhou et al. (2022b; a); Hao et al. (2024), LoRA Hu et al. (2021); Zhang et al. (2022); Xiong et al. (2024), and Adapter Houlsby et al. (2019); He et al. (2021); Chen et al. (2022); Xu et al. (2024b); Hao et al. (2023b; a), freeze pre-trained model parameters while adding extra tunable parameters to minimize storage demands. For parameter-efficient text-video retrieval, VoP Huang et al. (2023) and DGL Yang et al. (2024) introduce extra modules to generate prompt tokens and capture global video information. Despite their good performance with small tunable parameters, they face high inference costs due to the significant token number in videos.
@@
Token compression methods Ren et al. (2023); Liang et al. (2022); Wang et al. (2022a); Long et al. (2023); Rao et al. (2021); Kong et al. (2022); Chen et al. (2023a); Bolya et al. (2022); Marin et al. (2021) aim to reduce computational burden by reducing tokens while preserving the essential information.
Image token compression methods focus on pruning or merging tokens in a single image. DeCo Yao et al. (2024) adopts 2D Adaptive Pooling to downsample vision tokens at the spatial level for MLLMs. EVIT Liang et al. (2022) identifies the attentive tokens and fuses the inattentive tokens. DiffRate Chen et al. (2023a) performs token pruning and merging simultaneously. ToMe Bolya et al. (2022) reduces tokens in a transformer gradually by merging a fixed number of tokens in each block. Although ToMe has been applied to video classification by simultaneously processing all frame tokens in Spatiotemporal MAE Feichtenhofer et al. (2022), this implementation is incompatible with CLIP’s image encoder for text-video retrieval. When processing each sampled frame via CLIP’s image encoder, image token compression methods effectively reduce spatial redundancy by merging similar tokens within a single frame. However, they do not address the temporal redundancy across frames, which greatly contributes to computational overhead.
Video token compression methods Wang et al. (2022a); Ding et al. (2023); Ren et al. (2023) focus on pruning or merging tokens across frames. STA Ding et al. (2023) considers temporal redundancy and semantic importance, which is tailored for video architectures where all frame tokens are processed jointly. TESTA Ren et al. (2023) introduces temporal aggregation for video-language pre-training. We re-implement these video token compression methods on CLIP for text-video retrieval.
In this work, we focus on text-video retrieval using CLIP, where each sampled frame is processed as an independent token set. Existing token compression methods are limited to pruning or merging tokens within a single token set for an image or video, without addressing token compression across multiple sets or incorporating temporal fine-tuning. In contrast, we have explored a practical and feasible path to reach both superior performance and computational efficiency. By fruitfully integrating parameter-efficient fine-tuning and token compression techniques, we propose TempMe and reach state-of-the-art performance. TempMe can progressively merge different frame token sets, and thus minimize spatio-temporal redundancy and enhance temporal modeling across frames.
@@ 3.1 Preliminaries
Text-video retrieval involves leveraging textual queries to accurately retrieve related videos, as well as utilizing video queries to find associated textual descriptions. The primary goal is to bridge the semantic gap between textual description and video content by learning a similarity function $s(t,v)$. Here, $t$ denotes a text, and $v$ represents a sequence of sampled frame $\left\{I_{i}\right\}_{i=1}^{F}$, where $F$ is the number of sampled frames in time. We employ LoRA Hu et al. (2021) to adapt CLIP Radford et al. (2021) for text-video retrieval tasks. CLIP employs transformer that includes both multi-head self-attention (MHSA) and feed-forward network (FFN). The trainable matrices of LoRA can be merged, introducing no extra inference latency. Further details are provided in Appendix A.
Notably, the classic method, VoP Huang et al. (2023), does not implement specific designs for the text modality. It merely introduces prompts within the text encoder. Similar to VoP, we apply LoRA to the text encoder, while primarily concentrating on the video modality.
@@ 3.2 Temporal Token Merging
For efficient text-video retrieval, we freeze the pre-trained CLIP and merely train LoRA in both the image and text encoders. In this subsection, we focus on merging temporal tokens across multiple sampled frames to minimize temporal redundancy and extract more comprehensive video features.
@@
In text-video retrieval, the significant impact of temporal redundancy on model complexity has not been adequately explored. To mitigate temporal redundancy, we proposed the Progressive Multi-Granularity (PMG) framework that efficiently merges similar temporal tokens across different frames. In our PMG framework (see Figure 2), we introduce a two-stage merging approach: image merging and clip merging. Initially, ImgMe block independently encodes each single frame, merging similar tokens within the same frame and leveraging pre-trained image knowledge to capture low-level spatial details. Next, ClipMe block (see Figure 3) is proposed to aggregate short-frame clips into extended-frame clips, merging temporal tokens across different frames and learning unified and discriminative video features.
Specifically, each sampled frame $I$ can be regarded as a discrete video clip lasting only one frame $v^{f=1}$. Concatenating these consecutive clips yields an extended clip with increased temporal length. In the toy example presented in Figure 2, 8 individual frames $\left\{v^{f=1}_{i}\right\}_{i=1}^{8}$ are aggregated into four clips, each containing with two frames $\left\{v^{f=2}_{i}\right\}_{i=1}^{4}$. These clips are then aggregated to form a complete video clip with all 8 frames $\left\{v^{f=8}_{i}\right\}_{i=1}^{1}$. As clips are progressively aggregated, we merge a significant number of similar tokens that arise from overlapping information across frames. Simultaneously, comprehensive video information is captured through the attention mechanism that processes tokens across different frames.
@@
Inspired by ToMe Bolya et al. (2022), the image merging step merges tokens to reduce by $r$ per layer, which is employed between the MHSA and FFN branches of each ImgMe Block. Following Bipartite Soft Matching Bolya et al. (2022), we first partition the tokens into two sets $\mathbb{A}$ and $\mathbb{B}$ by alternating. For each token in set $\mathbb{A}$, we find its most similar token in Set $\mathbb{B}$ using cosine similarity. The top $r$ most similar pairs are then merged by averaging their features. Given an input $\mathbb{R}^{1\times N\times D}$, the image merging stage with $K$ ImgMe Block outputs $\mathbb{R}^{1\times(N-rK)\times D}$, where $N$ denotes the number of tokens, and $D$ denotes the dimension of the token feature.
To prevent loss of image information, $r$ is set to a low value: $r=2$ for CLIP-ViT-B/32 and $r=10$ for CLIP-ViT-B/16. However, the model complexity remains high due to the processing of multiple sampled frames. Therefore, we aim to merge temporal tokens to further reduce model complexity.
@@
Instead of ImgMe Block which merges tokens in each single frame, we propose ClipMe Block to process multi-frame clips as shown in Figure 3. In ClipMe Block, we introduce two novel steps: cross-clip merging and intra-clip merging. During cross-clip merging, adjacent clips are aggregated, which significantly reduces the number of temporal tokens and generates a new clip. The intra-clip merging step further compresses the tokens within the newly formed clip.
In cross-clip merging, we aim to combine tokens with their most similar tokens from the same subject, whether these matches occur within a single frame or span across multiple frames. Initially, all tokens from the input $f$ clips $\mathbb{R}^{f\times N\times D}$ are alternately divided into Set $\mathbb{A}$ or Set $\mathbb{B}$. This alternating assignment scheme is based on the observation that the adjacent tokens often depict the same subject, thus their features tend to be similar. Moreover, this assignment ensures that tokens representing the same subject across various frames are distributed between both sets. Considering the high temporal redundancy observed in videos, we select a large subset, $fN\times(1-R_{c})$, of the most similar token pairs for merging, where $R_{C}$ denotes the ratio of kept tokens in cross-clip merging. Before merging, the trainable clip positional embeddings are added to facilitate temporal understanding in MHSA. Subsequently, these merged tokens, along with the unmerged tokens, constitute a new clip $\mathbb{R}^{1\times(fN\times R_{c})\times D}$, which is then fed into MHSA.
During intra-clip merging, the tokens in the newly formed clip, improved by MHSA for better temporal and spatial details, are further reduced. In contrast to the image merging step that merges a small number $r$ of tokens in an individual frame, the intra-clip merging step merges a larger proportion of tokens within a clip containing multiple frames, compressing $\mathbb{R}^{1\times N\times D}$ into $\mathbb{R}^{1\times(N\times R_{I})\times D}$, where $R_{I}$ denotes the ratio of kept tokens in intra-clip merging.
Table 1: Complexity comparisons. We report the R-Sum metric for 12 sampled frames on MSRVTT and 64 sampled frames on ActivityNet, respectively. The best metric is highlighted in bold.
@@
Method GFLOPs # Tokens R@1/R-Sum GFLOPs # Tokens R@1/R-Sum GFLOPs # Tokens R@1/R-Sum
LoRA 53.0 (100%) $12\times 50$ (100%) 43.7/193.0 276.7 (100%) $64\times~{}~{}50$ (100%) 38.7/191.5 211.3 (100%) $12\times 197$ (100%) 47.3/201.4
DiffRate 36.8 (  69%) $12\times 20$ (  40%) 41.5/189.9 190.1 (  69%) $64\times~{}~{}20$ (  40%) 38.0/188.7 138.5 (  66%) $12\times~{}~{}49$ (  25%) 47.3/202.4
ToMe 40.2 (  76%) $12\times 26$ (  52%) 42.9/191.4 208.5 (  75%) $64\times~{}~{}26$ (  52%) 38.6/189.6 144.4 (  68%) $12\times~{}~{}77$ (  39%) 46.2/200.6
TempMe 34.8 (  65%) $\mathbf{1\times 97}$ (  16%) 46.1/198.6 180.3 (  65%) $\mathbf{1\times 500}$ (  16%) 44.9/205.6 121.4 (  57%) $\mathbf{1\times 127}$ (    5%) 49.0/206.7